# Relative Representations: Topological and Geometric Perspectives

### v2

@@ 2.2 Symmetry Groups of Activation Functions
$\varphi_{m}(x,\widetilde{W})$ $=\lambda_{\sigma}(A_{m})\circ\varphi_{m}(x,W),$
$\gamma_{m}(x,\widetilde{W})$ $=\gamma_{m}(x,W)\circ\lambda_{\sigma}(A_{m})^{-1}.$
@@ 2.3 Topological Densification
Connected components of the truncation graph can be interpreted as clusters of data. This is the idea behind density-based spatial clustering methods [14]. Instead, hierarchical density-based clustering [6] – as well as the related notion of 00-th persistent homology – considers the evolution of clusters as the parameter $\varepsilon$ varies. For $\varepsilon=0$ all the points in $\mathcal{D}$ form a distinct connected component. As $\varepsilon$ grows, edges are added to the truncation graph. This implies that if $\varepsilon\leq\delta$, each component in $\Gamma_{\varepsilon}(\mathcal{D})$ is contained in a component of $\Gamma_{\delta}(\mathcal{D})$. Moreover, two distinct components in $\Gamma_{\varepsilon}(\mathcal{D})$ can merge together and become connected in $\Gamma_{\delta}(\mathcal{D})$. The value of $\varepsilon$ at which two connected components merge is called a death time. Since more than two connected components can merge at the same parameter value, death times form a multi-set in $\mathbb{R}_{\geq 0}$ with multiplicity given by the number of merged components minus one. This multi-set is denoted by $\dagger(\mathcal{D})$, and it is equivalent to the persistence diagram of 00-dimensional persistent homology. It can be shown that death times coincide exactly with the lengths of the edges of the minimum spanning tree of $\mathcal{D}$, i.e., the connected tree with vertices $\mathcal{D}$ and that is of minimal total edge length. This enables to compute the persistence diagram via, for example, Kruskal’s algorithm [26], whose time complexity is $\mathcal{O}(|\mathcal{D}|^{2}\log|\mathcal{D}|)$.
@@
We remark that the content of this section can be extended to higher-dimensional persistent homology [13]. The above definitions corresponds to an explicit construction of the 00-dimensional persistent homology of the Vietoris-Rips complex, given that in this case all the birth times are 00. By considering $i$-th dimensional homology, it is possible to define persistence diagrams for every $i\geq 0$ – see Appendix (Section A). The topological loss can then be extended based on features of the points in the corresponding persistence diagrams. Moreover, other filtration constructions, such as the more efficient Witness complexes [36], can replace the Vietoris Rips complex. However, following [21], we focus on the 00-dimensional Vietoris-Rips persistent homology for simplicity, leaving further investigation of higher-dimensional extensions and other filtrations for future work.
@@ 4 Experiments
We perform a random subsampling of 1% of the Amazon Reviews dataset [24], where the task consists in predicting user ratings in a range from $1$ to $5$ stars. We consider such task in two languages: English (‘en’) and French (‘fr’). The anchor sets $\mathcal{A}$ and $\mathcal{A}^{*}$ (notation from Section 2.1) consist of translated texts. Specifically, we first select a number of anchors from the English dataset equal to the dimensionality of the latent space ($768$ dimensions), and then translate them to other languages using Google Translate. A Python code implementing our models and experiments is available at a public anonymized repository: https://tinyurl.com/TopoRelTrans.
@@ Appendix A Extended Topological Densification
In Section 2.3 we describe 00-th persistent homology in relation to topological densification [21]. We now present a concise definition of $n$-th persistent homology for $n\geq 0$, and elaborate on possible generalizations of the topological densification loss (Definition 2.4). For an introduction to persistent homology and topological data analysis we refer the reader to [8].
@@
From Geometry to Algebra. The topology of a simplicial complex and in particular its homology can be computed in an algorithmic way through simplicial homology. Fixed a natural number $n$ and a coefficient field, the $n$-th homology of a simplicial complex is a vector space encoding connectivity of the simplicial complex. The 00-th homology describes connected components, $1$-st homology describes cycles that are not bounded by faces, $2$-nd homology describes voids and for $n\geq 2$ one can think of higher dimensional analogues. In the cases of a sequence of simplicial complexes, $n$-th homology can be applied to the whole sequence to obtain a sequence of vector spaces and linear maps. This is referred to $n$-th persistent homology module.
@@
where $\dagger(\varphi(\mathcal{D}_{i}))$ is the set of death times in the persistence diagram of $\varphi(D_{i})$. For 00-th persistence, death times correspond to lengths of bars in a barcode, since all bars start at parameter value 00. Consider now the case of $n$-th persistent homology of the $\varphi(\mathcal{D}_{i})$ and the corresponding persistence diagrams $\ddagger(\varphi(\mathcal{D}_{i}))$. The formula for of the topological densification loss can be generalized as:
@@ Appendix B Batch Construction
    For each datapoint $x$ in the batch, a sub-batch is constructed by sampling with replacement $n-1$ more datapoints with the same class as $x$.

### v3

@@ 2.2 Symmetry Groups of Activation Functions
[…]
@@ 2.3 Topological Densification
Connected components of the truncation graph can be interpreted as clusters of data. This is the idea behind density-based spatial clustering methods [14]. Instead, hierarchical density-based clustering [6] – as well as the related notion of 0-th persistent homology – considers the evolution of clusters as the parameter $\varepsilon$ varies. For $\varepsilon=0$ all the points in $\mathcal{D}$ form a distinct connected component. As $\varepsilon$ grows, edges are added to the truncation graph. This implies that if $\varepsilon\leq\delta$, each component in $\Gamma_{\varepsilon}(\mathcal{D})$ is contained in a component of $\Gamma_{\delta}(\mathcal{D})$. Moreover, two distinct components in $\Gamma_{\varepsilon}(\mathcal{D})$ can merge together and become connected in $\Gamma_{\delta}(\mathcal{D})$. The value of $\varepsilon$ at which two connected components merge is called a death time. Since more than two connected components can merge at the same parameter value, death times form a multi-set in $\mathbb{R}_{\geq 0}$ with multiplicity given by the number of merged components minus one. This multi-set is denoted by $\dagger(\mathcal{D})$, and it is equivalent to the persistence diagram of 0-dimensional persistent homology. It can be shown that death times coincide exactly with the lengths of the edges of the minimum spanning tree of $\mathcal{D}$, i.e., the connected tree with vertices $\mathcal{D}$ and that is of minimal total edge length. This enables to compute the persistence diagram via, for example, Kruskal’s algorithm [26], whose time complexity is $\mathcal{O}(|\mathcal{D}|^{2}\log|\mathcal{D}|)$.
@@
We remark that the content of this section can be extended to higher-dimensional persistent homology [13]. The above definitions corresponds to an explicit construction of the 0-dimensional persistent homology of the Vietoris-Rips complex, given that in this case all the birth times are 0. By considering $i$-th dimensional homology, it is possible to define persistence diagrams for every $i\geq 0$ – see Appendix (Section A). The topological loss can then be extended based on features of the points in the corresponding persistence diagrams. Moreover, other filtration constructions, such as the more efficient Witness complexes [36], can replace the Vietoris Rips complex. However, following [21], we focus on the 0-dimensional Vietoris-Rips persistent homology for simplicity, leaving further investigation of higher-dimensional extensions and other filtrations for future work.
@@ 4 Experiments
We perform a random subsampling of 1% of the Amazon Reviews dataset [24], where the task consists in predicting user ratings in a range from $1$ to $5$ stars. We consider such task in two languages: English (‘en’) and French (‘fr’). The anchor sets $\mathcal{A}$ and $\mathcal{A}^{*}$ (notation from Section 2.1) consist of translated texts. Specifically, we first select a number of anchors from the English dataset equal to the dimensionality of the latent space ($768$ dimensions), and then translate them to other languages using Google Translate. A Python code implementing our models and experiments is available at a public repository: https://github.com/AGarciaCast/TopoRobRelTrans.
@@ Appendix A Extended Topological Densification
In Section 2.3 we describe 0-th persistent homology in relation to topological densification [21]. We now present a concise definition of $n$-th persistent homology for $n\geq 0$, and elaborate on possible generalizations of the topological densification loss (Definition 2.4). For an introduction to persistent homology and topological data analysis we refer the reader to [8].
@@
From Geometry to Algebra. The topology of a simplicial complex and in particular its homology can be computed in an algorithmic way through simplicial homology. Fixed a natural number $n$ and a coefficient field, the $n$-th homology of a simplicial complex is a vector space encoding connectivity of the simplicial complex. The 0-th homology describes connected components, $1$-st homology describes cycles that are not bounded by faces, $2$-nd homology describes voids and for $n\geq 2$ one can think of higher dimensional analogues. In the cases of a sequence of simplicial complexes, $n$-th homology can be applied to the whole sequence to obtain a sequence of vector spaces and linear maps. This is referred to $n$-th persistent homology module.
@@
where $\dagger(\varphi(\mathcal{D}_{i}))$ is the set of death times in the persistence diagram of $\varphi(D_{i})$. For 0-th persistence, death times correspond to lengths of bars in a barcode, since all bars start at parameter value 0. Consider now the case of $n$-th persistent homology of the $\varphi(\mathcal{D}_{i})$ and the corresponding persistence diagrams $\ddagger(\varphi(\mathcal{D}_{i}))$. The formula for of the topological densification loss can be generalized as:
@@ Appendix B Batch Construction
    For each datapoint $x$ in the batch, a sub-batch is constructed by sampling with replacement n1n-1 more datapoints with the same class as $x$.